OpenStudy (swag):

describe how you would write the equation of a line given a point on the line and a parallel line.

OpenStudy (swag):

@tafkas77

OpenStudy (anonymous):

Whoops, sorry SWAG. I logged off, but saw your mention in my email. Let me take a look at this.

OpenStudy (swag):

Haha its ok I feel special that you came to help haha

OpenStudy (swag):

also the value of f[g(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5

OpenStudy (anonymous):

Ha, thanks. That's what I do! :) well, first things first. do you know anything about parallel lines or are you kind of looking at these questions like, 0_o

OpenStudy (swag):

O.o lolol

OpenStudy (anonymous):

Ha ha, okay! Well let me get you started then :D

OpenStudy (anonymous):

Lines that are parallel have the same slope. If you know the slope of the parallel line, you also know the slope of the line that you're trying to find. So for example, if you were given a line with a slope of 1/2 (such as y=1/2x), and you needed a parallel line, you could write an equation like this: y=1/2x + 2. Do you get this?

OpenStudy (swag):

Kind of yea

OpenStudy (anonymous):

Okay. Let me know if you're still confused, otherwise I'll move on to your next question.

OpenStudy (swag):

I got it

OpenStudy (anonymous):

Great. Moving on; it might take a second for me to type this up; okay? Do you need the whole thing; is this just so you can understand how to do it? You're completely lost?

OpenStudy (swag):

No this question i just need the steps and awnser

OpenStudy (anonymous):

Cool. Let me type it up for ya. :)

OpenStudy (swag):

k

OpenStudy (anonymous):

"What is the value of f[g(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?" First, you have to take the equation g(x) = 2x - 5 and plug it into the equation f(x)= 2x + 1 for the x value (there's only one in the f(x) equation, so this problem isn't too bad): f[g(x)] = 2(2x - 5) + 1 Now, you have to plug in 4 for EVERY x value in the ENTIRE equation: f[g(4)] = 2[2(4) - 5] + 1 and simplify. f[g(4)] = 2(8 - 5) + 1 f[g(4)] = 2(3) + 1 f[g(4)] = 6 + 1 f[g(4)] = 7 That is your answer. Do you understand how I did it?

OpenStudy (swag):

Yes, you explained it a whole lot better than my teacher did. your a life saver

OpenStudy (anonymous):

Really? Wow! Thanks! I'm glad I could help. :)